A carpenter charges a set fee for each job in additon to his hourly charge. If a carpenter bills a customer $150 for a job which takes him 3 hours and bills a second customer $180 for a 6 hour job, how much would he charge for a job that takes him 9 hours?
Found 2 solutions by stanbon, KissedPeach:
Answer by stanbon(48568)

A carpenter charges a set fee for each job in additon to his hourly charge. If a carpenter bills a customer $150 for a job which takes him 3 hours and bills a second customer $180 for a 6 hour job, how much would he charge for a job that takes him 9 hours?
------------------
You have two points relating time and charge: (3,150) and (6,180)
slope = 30/3 = 10
150 = 10*3 + b
b = 120
EQUATION:
charge = 10(hrs)+120
C(9) = 10*9 + 120
C(9) = $210 (charge for a 9 hour job)
================
Cheers,
Stan H.

Answer by KissedPeach(1)

Using the five step method your answer would be.
1. The problem asks for the amount charged for a 9 hour job.
2. let x= the set fee and let y= the hourly charge.
3 (150-3y) + 6y = 180
4 (150-3y)+6y=180
150-150 -3y + 6y =180
-3y+6y = 30
3y=30
y=10
x=150-(3*10)
x=150-30
x=120

120+(9*10)=x
120+90=x
210=x
5. 120+(3*10)=150
120+30=150
150=150
120+(6*10)=180
120+60=180
180=180
120+(9*10)=210
120+10=210
210=210
the carpenter charges $210 for a 9 hour job.